OpenStudy (goose.):

Multiply (4x + 7)^2. 16x^2 + 49 16x^2 + 28x + 49 16x^2 + 56x + 49 8x^2 + 14

OpenStudy (texaschic101):

(4x + 7) ^2 = (4x + 7)(4x + 7) do you know the FOIL method ?

OpenStudy (goose.):

To solve this I would do (4x + 7)(4x + 7) Correct?

OpenStudy (texaschic101):

yes

OpenStudy (goose.):

I know how to do it, kind of.. Walk em through it?

OpenStudy (goose.):

me*

OpenStudy (goose.):

It would be 4x times 4x then 4x times 7?

OpenStudy (texaschic101):

(4x + 7)(4x + 7) using FOIL -- first,outside,inside,last) First -- multiply first terms in each set....4x * 4x = 16x^2 Outside -- multiply outside terms in each set...4x * 7 = 28x Inside -- multiply inside terms in each set...7 * 4x = 28x Last -- multiply last terms in each set...7 * 7 = 49 put it all together....16x^2 + 28x + 28x + 49 -- combine like terms 16x^2 + 56x + 49 questions ?

OpenStudy (goose.):

Nope, I got it, thank you!

OpenStudy (texaschic101):

anytime :)